OpenStudy (anonymous):

If 1 + 1 = 2 then what is the circumference of the sun?

OpenStudy (anonymous):

4.4 million km

OpenStudy (anonymous):

99% chance that you googled that. but thanks anyways

OpenStudy (anonymous):

Sigh... http://24.media.tumblr.com/tumblr_mbt2weCL1g1rtrx1ho1_400.gif

OpenStudy (nincompoop):

we usually use the sun's diameter or radius and distance from the earth or from the moon like ALL the forking time so you can shove google up your piehole

OpenStudy (anonymous):

I've taken like 50 astronomy classes, and accuses me of google... C = 2*pi*r r = 675km = 4.24 million km, slightly off, but the numbers are also an approximate.

OpenStudy (anonymous):

And yes, in the cosmos 160,000 km is nothing.

OpenStudy (anonymous):

Okkkkkk nin relax, this guy's asking 1+1 = 2 and you take out the integration..

OpenStudy (nincompoop):

we can also show him how we are able to estimate the size of the sun by mere trig